# Structured dataset of reported cloud seeding activities in the United States (2000–2025) using a large language model

### v4

@@ Abstract
Cloud seeding, a weather modification technique used to increase precipitation, has been employed in the western United States since the 1940s. However, structured datasets are not currently available to analyze these efforts. To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S. from 2000–2025, including the project name, year, season, state, operator, seeding agent, apparatus used for deployment, stated purpose, target area, control area, start date, and end date. Using OpenAI’s o3 large language model (LLM), combined with our multi-stage PDF-to-text extraction pipeline, we processed 832 historical reports from the National Oceanic and Atmospheric Administration (NOAA). The resulting dataset demonstrates 98.38% estimated accuracy, based on manual review of a random sample of 200 records, and is publicly available on Zenodo. Our results help fill the gap in structured cloud seeding data and demonstrate the potential for LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
To solve this data access problem, we created a structured dataset from 832 NOAA cloud seeding reports spanning from 2000 to 2025. Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S. weather modification activities over time. The dataset achieves an estimated 98.38% accuracy across all extracted fields, based on a manual evaluation of 200 randomly sampled records (out of 832), each reviewed by two independent human annotators against the original PDFs.
This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting. By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest[4].
@@
For example, while there have been several recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g. the 2012 ASCII (AgI Seeding Cloud Impact Investigation) over Sierra Madre range in Wyoming or the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[5, 6]), there has been little qualitative analysis of reported cloud seeding activities. With recent renewed interest in both traditional weather modification technology (e.g. the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geoengineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[7], this dataset can provide historical context for weather modification usage in the US. In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification[4].
@@ Methods
Figure 1: NOAA Data Extraction Pipeline
@@
Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports Archive [3] (https://library.noaa.gov/weather-climate/weather-modification-project-reports). A total of 1,025 PDF files were collected, including Form 17-4, Form 17-4A, and other supplemental documents. These files represent all publicly reported weather modification activities submitted to NOAA in the United States from 2000 to 2025. The reports are regulatory submissions intended for public access and are in the public domain. Their reuse and redistribution are permitted under U.S. government policy.
Each PDF included the initial Form 17-4, a follow-up Form 17-4A, or both. Form 17-4 contains project-level summaries, including the purpose of the project, the planned operational period, and the seeding agents and apparatus used. Form 17-4A provides more detailed tabular information on seeding activity, such as the number of days seeded per month and the quantities of agent released.
For this dataset, we focused exclusively on extracting fields from Form 17-4, which contained the majority of the relevant descriptive information. While many PDFs also included supporting materials such as maps, meteorological analyses, operational logs, and narrative summaries, these varied widely in structure and content. As a result, they were excluded from this version of the dataset due to their inconsistent formatting and the additional preprocessing they would require.
@@
A custom Python pipeline was developed to organize and preprocess all Form 17-4 PDF documents (Figure 1). First, all downloaded files were stored in a local directory. In many cases, a single cloud seeding project was documented across multiple submissions, indicated by suffixes such as .M (initial report), .I (interim report), and .F (final report). Since these reports refer to the same geographic area and operational period, they were programmatically consolidated to preserve project-level information and ensure completeness during data extraction. After consolidation, there were 832 unique weather modification projects to process and include in our dataset.
The scanned documents included both typed and hand-written text (and sometimes a mix of both in the same document). To convert the PDFs into plain text for further processing, a combination of three technologies was used. For native, text-based documents (57% of documents), text was extracted using the pymupdf Python library [8]. For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [9] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology that combines OCR and native-text extraction [10]. All code for the preprocessing pipeline, including file merging and text extraction, is available in a public GitHub repository [11].
@@
We experimented with multiple prompt templates using prompt engineering techniques to improve extraction quality. The final prompt, shown below with truncation, employs chain-of-thought reasoning by guiding the model through intermediate steps before producing structured outputs. This reasoning style improves performance on tasks requiring multi-step inference and disambiguation [13]. The prompt was used with the o3 model for all extractions in the final dataset. All prompt versions and evaluation scripts are available in our GitHub repository [11].
@@
As instructed, the LLM analyzed the extracted PDF text and responded with a list containing the key information we requested. In order to neatly store this information into CSV format, we then mapped key values into more concise column names, stripped any blank space, and converted to lowercase. Using the formatted data object, each project could be easily written to CSV, with one row per object and columns corresponding to the extracted fields. Lastly, to prepare the final dataset, we used clean-dataset.py [11] to remove any duplicates, standardize formatting, and fill blank cells with NaN.
@@
The resulting dataset is publicly available on Zenodo [14] as a single CSV file. It contains structured information from 832 unique cloud seeding project reports across the United States from 2000 to 2025, with one row per project and one column per extracted field (Figure 2).
Figure 2: Dataset Screenshot (cloud_seeding_us_2000_2025.csv)
@@
Figure 3: U.S. Cloud Seeding Activity by State (2000–2025). States with active weather modification programs show the highest number of recorded operations. Specific locations (shown as salmon markers) are from geocoding target_area and state fields with the GoogleMaps API and have not been verified for accuracy.
Figure 4: Cloud Seeding Activity by State over Time (2000–2025). The number of activities peaked around 2005, declined gradually, and rose again after 2023.
Figure 5: Purpose of Cloud Seeding Activity (2000–2025). Snowpack enhancement is the leading purpose, followed by increasing precipitation and rain.
Figure 6: Agent and Apparatus Used for Cloud Seeding (2000–2025). Silver iodide dominates among agents and is deployed most frequently using ground-based apparatus.
@@
To address the potential for LLMs to hallucinate content not present in the source documents, and the inconsistent quality of the input PDFs themselves, we measured field-level extraction accuracy by manually reviewing a random sample of 200 records extracted by the LLM and comparing them to ground truth.
@@
Accuracy Evaluations
From the population of N=832 records we randomly sampled n=200 to evaluate field-level extraction accuracy (Table 1). This evaluation was conducted by two independent human experts manually creating a golden dataset (ground truth) using the source PDFs, then comparing the extracted fields against the golden dataset using compare-to-golden.py [11]. Assuming a true population accuracy of 95%, this sample size was chosen to yield a margin of error of $\pm 3$ percentage points at the 95% confidence level.
@@
The o3 reasoning model achieved the highest overall accuracy at 96.33%, but was initially an order of magnitude more expensive than o4-mini costing approximately \$0.05 per document versus \$0.005 for o4-mini, which achieved 95.00% accuracy. However, in June 2025, OpenAI reduced the price of o3, bringing its cost down to \$0.01 per document. Following this price drop, we adopted o3 to maximize accuracy without increasing cost.
Table 2: Model Comparison (n=50; prompt C)
LLM Overall Average Accuracy (%)
gpt-4.1 93.33%
gpt-4.1-mini 93.33%
gpt-4o-mini 91.67%
o3 98.38%
o4-mini 95.00%
@@
Table 3: Prompt Comparison (n=50) using o4-mini LLM
Prompt Overall Average Accuracy (%)
Prompt A (Detailed) 93.67%
Prompt B (Concise) 44.00%
Prompt C (Chain-of-Thought) 95.00%
Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering. In particular, chain-of-thought reasoning and example-driven prompts are well-suited for scientific data extraction from complex historical documents.
@@
Python code for extracting the dataset from the NOAA reports can be obtained from the public repository [11]. Python code for analyzing and visualizing the dataset is also publicly available on GitHub [17].

### v5

# Structured dataset of reported cloud seeding activities in the United States (2000–2025) using an LLM

@@ Abstract
Cloud seeding, a weather modification technique used to increase precipitation, has been practiced in the western United States since the 1940s. However, comprehensive datasets are not currently available to analyze these efforts. To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S. from 2000–2025, including the project name, year, season, state, operator, seeding agent, apparatus used for deployment, stated purpose, target area, control area, start date, and end date. Combining our multi-stage PDF-to-text extraction pipeline with OpenAI’s o3 large language model (LLM), we processed 832 historical reports from the National Oceanic and Atmospheric Administration (NOAA). The resulting dataset demonstrates 98.38% estimated accuracy, based on manual review of 200 randomly sampled records, and is publicly available on Zenodo. This dataset addresses the gap in cloud seeding data and demonstrates the potential for LLMs to extract structured information from historical environmental documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
To solve this data access problem, we created a structured dataset from 832 NOAA cloud seeding reports spanning from 2000 to 2025. Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S. weather modification activities over time. The dataset achieves an estimated 98.38% accuracy across all extracted fields, based on a manual evaluation of 200 randomly sampled records (out of the total population of 832), each reviewed by two independent human annotators against the original PDFs.
This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management, many of which exist in PDF form with highly variable formatting. By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest[4].
@@
For example, while there have been several recent research campaigns focused on process-level understanding and evaluation of cloud seeding, such as the the 2012 ASCII (AgI Seeding Cloud Impact Investigation) over the Sierra Madre range in Wyoming, or the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[5, 6], there has been little quantitative analysis of reported cloud seeding activities. With recent renewed interest in both traditional weather modification technology (e.g. the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as geoengineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[7], this dataset can provide historical context for weather modification usage in the United States. In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification[4].
@@ Methods
Figure 1: NOAA Form 17-4 Dataset Extraction Pipeline
@@
Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports Archive (https://library.noaa.gov/weather-climate/weather-modification-project-reports) [3]. A total of 1,025 PDF files were collected, including Form 17-4, Form 17-4A, and other supplemental documents. These 1,025 files represent all publicly reported weather modification activities submitted to NOAA. Since the reports are regulatory submissions intended for public access, they are in the public domain and their reuse and redistribution are permitted under U.S. government policy.
Each PDF included Form 17-4, Form 17-4A, or both, along with supplemental pages. Form 17-4 contains the main project summary, including the project purpose, operational period, operator, seeding agents, apparatus for deployment, and target/control areas. Form 17-4A provides tabular data on seeding activity, including days seeded per month and agent quantities released.
For our dataset, we focused exclusively on extracting fields from Form 17-4 which contained the majority of relevant project descriptive information. While many PDFs included supporting materials, such as maps, meteorological analyses, operational logs, and narrative summaries, these were excluded due to inconsistent formatting that would require extensive preprocessing with minimal dataset enhancement.
@@
A custom Python pipeline was developed to preprocess all Form 17-4 PDF documents (Figure 1). First, all downloaded files were stored in a local directory. In many cases, a single cloud seeding project was documented across multiple submissions, indicated by suffixes such as .M (initial report), .I (interim report), and .F (final report). These reports were consolidated into a single file per project using a combination of programmatic and manual methods to preserve project-level information. After consolidation, 832 unique weather modification projects remained for processing and inclusion in the dataset.
The scanned documents included both typed and hand-written text, and sometimes a mix of both in the same document. To convert the PDFs into plain text for further processing, a combination of three technologies was used. For native, text-based documents (57% of documents), text was extracted using the pymupdf Python library [8]. For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [9] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology that combines OCR and native-text extraction [10]. All code for the preprocessing pipeline, including file merging, text extraction, and LLM-integration, is publicly available on GitHub (https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor) [11].
@@
We experimented with multiple prompt templates using prompt engineering techniques to improve extraction quality. The final prompt, shown below with truncation, employs chain-of-thought reasoning by guiding the model through intermediate steps before producing structured outputs. This reasoning style improves performance on tasks requiring multi-step inference and disambiguation [13]. The prompt was used with the o3 reasoning model for all extractions in the final dataset. All prompt versions and evaluation scripts are available in our public GitHub repository [11].
@@
As instructed, the LLM analyzed the extracted PDF text and responded with a list containing the key information we requested. In order to neatly store this information into CSV format, we then mapped key values into more concise column names, stripped any blank space, and converted to lowercase. Using the formatted data object, each project could be easily written to CSV, with one row per object and columns corresponding to the extracted fields. Lastly, to prepare the final dataset, we used clean-dataset.py in [11] to remove any duplicates, standardize formatting, and fill blank cells with NaN.
@@
The dataset is available on Zenodo as a single CSV file (https://doi.org/10.5281/zenodo.14925811) [14]. It contains structured information from 832 unique cloud seeding project reports across the United States from 2000 to 2025, with one row per project and one column per extracted field (Figure 2).
Figure 2: Screenshot of the CSV Dataset (cloud_seeding_us_2000_2025.csv)
@@
Figure 3: Cloud Seeding Activity by U.S. State (2000–2025). States with active weather modification programs show the highest number of recorded operations. Specific locations (shown as salmon markers) are from geocoding the stated target_area field with the GoogleMaps API.
Figure 4: Cloud Seeding Activity by U.S. State over Time (2000–2025). The number of activities peaked between 2003-2005, declined gradually, and rose again after 2021.
Figure 5: Stated Purpose of Cloud Seeding Activity (2000–2025). Augmenting snowpack is the leading purpose, followed by increasing precipitation.
Figure 6: Agent and Apparatus Used for Cloud Seeding Operations (2000–2025). Silver iodide dominates among agents and is deployed most frequently using a ground-based apparatus.
@@
To address the potential for LLMs to hallucinate content not present in the source documents, and the inconsistent quality of the input PDFs themselves, we measured field-level extraction accuracy by manually reviewing a random sample of 200 records extracted by the LLM and comparing them to ground truth. This evaluation provides an empirical bound on potential extraction errors in the released dataset, enabling downstream users to incorporate measurement uncertainty into their analyses.
@@
Extraction Accuracy
From the population of N=832 records we randomly sampled n=200 to evaluate field-level extraction accuracy (Table 1). This sample size was chosen to yield a margin of error of $\pm 3$ percentage points at the 95% confidence level. The evaluation was conducted by two independent human experts manually creating a golden dataset (ground truth) using the source PDFs, then comparing the extracted fields against the golden dataset using compare-to-golden.py [11].
@@
The o3 reasoning model achieved the highest overall accuracy at 96.33%, but was initially an order of magnitude more expensive than o4-mini costing approximately \$0.05 per document versus \$0.005 for o4-mini, which achieved 95.00% accuracy. However, in June 2025, OpenAI reduced the price of o3, bringing its cost down to \$0.01 per document. Following this price drop, we adopted o3 to maximize accuracy while maintaining low cost.
Table 2: Extraction Accuracy by LLM (n=50; prompt C)
LLM Accuracy (%) 95% Confidence Interval
gpt-4.1 93.33% [86.11%, 97.48%]
gpt-4.1-mini 93.33% [86.11%, 97.48%]
gpt-4o-mini 91.67% [84.03%, 96.34%]
o3 96.33% [88.08%, 98.67%]
o4-mini 95.00% [88.08%, 98.67%]
@@
Table 3: Accuracy by LLM Prompt (n=50) using o4-mini
Prompt Accuracy (%) 95% Confidence Interval
Prompt A 93.67% [86.64%, 97.56%]
Prompt B 44.00% [30.98%, 57.69%]
Prompt C 95.00% [88.08%, 98.67%]
Together, these validation efforts show that LLM-based data extraction accuracy can be significantly improved through deliberate prompt design and model selection. In particular, prompts incorporating chain-of-thought reasoning and field-specific examples consistently enhanced extraction accuracy, even in complex historical documents with heterogeneous formatting and quality. These findings suggest that careful prompt engineering should be considered a core methodological step when using LLMs to extract scientific data from historical documents.
@@
Data Availability
The dataset is publicly available on Zenodo (https://doi.org/10.5281/zenodo.14925811) [14]. The original NOAA files used as input are publicly available on NOAA (https://library.noaa.gov/weather-climate/weather-modification-project-reports) [3].
@@
Python code for extracting the dataset from the NOAA reports can be obtained from our public GitHub repository (https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor) [11]. Python code for visualizing the dataset is also publicly available on GitHub (https://github.com/jdonohue44/US-Cloud-Seeding-Analysis) [17].